Let f(x) = x^a, xER^n, where a is a multi-index. Compute f-hat.
1
Expert's answer
2019-10-28T18:40:18-0400

"\hat{f}(\omega)=\frac{1}{(2\pi)^{\frac{n}{2}}}\int\limits_{\mathbb R^n}f(x)e^{-i(x,\omega)}dx"

"\hat{f}(\omega)=\frac{1}{(2\pi)^{\frac{n}{2}}}\int\limits_{\mathbb R^n}x^{\alpha}e^{-i(x,\omega)}dx=\frac{1}{(2\pi)^{\frac{n}{2}}}\int\limits_{\mathbb R^n}\prod\limits_{k=1}^n x_k^{\alpha_k}e^{-ix_k\omega_k}dx="

"=\prod\limits_{k=1}^n\bigl(\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x_k^{\alpha_k}e^{-ix_k\omega_k}dx_k\bigr)"

Consider "\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^{\alpha}e^{-ix\omega}dx". Prove by induction that "\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^{\alpha}e^{-ix\omega}dx=i^{\alpha}\sqrt{2\pi}\delta^{(\alpha)}(\omega)".

Since "\delta(\omega)=\overline{\delta(\omega)}" , for "\alpha=0" we have "\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}e^{-ix\omega}dx=\overline{\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}e^{ix\omega}dx}="

"=\sqrt{2\pi}\cdot\overline{\delta(\omega)}=\sqrt{2\pi}\delta(\omega)".

Suppose that "\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^ke^{-ix\omega}dx=i^k\sqrt{2\pi}\delta^{(k)}(\omega)", where "k\ge 0".

Then "i^{k+1}\sqrt{2\pi}\delta^{(k+1)}(\omega)=i\frac{d}{d\omega}\bigl(i^k\sqrt{2\pi}\delta^{(k)}(\omega)\bigr)". By the induction hypothesis we obtain "i^{k+1}\sqrt{2\pi}\delta^{(k+1)}(\omega)=i\frac{d}{d\omega}\bigl(\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^ke^{-ix\omega}dx)\bigr)="

"=\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}ix^k\frac{de^{-ix\omega}}{d\omega}dx=\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^{k+1}e^{-ix\omega}dx"

So we prove this statement for "\alpha=k+1".

By the principle of mathematical induction we obtain that "\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x^{\alpha}e^{-ix\omega}dx=i^{\alpha}\sqrt{2\pi}\delta^{(\alpha)}(\omega)" for all "\alpha\ge 0".

So "\hat{f}(\omega)=\prod\limits_{k=1}^n\bigl(\frac{1}{\sqrt{2\pi}}\int\limits_{\mathbb R}x_k^{\alpha_k}e^{-ix_k\omega_k}dx_k\bigr)="

"=\prod\limits_{k=1}^n\bigl(i^{\alpha_k}\sqrt{2\pi}\delta^{(\alpha_k)}(\omega_k)\bigr)=i^{|\alpha|}(2\pi)^\frac{n}{2}\prod\limits_{k=1}^n\delta^{(\alpha_k)}(\omega_k)"

Answer: "i^{|\alpha|}(2\pi)^\frac{n}{2}\prod\limits_{k=1}^n\delta^{(\alpha_k)}(\omega_k)"